On a map 1 inch equal 20 miles - two cities are 7 inches apart on a map - what is the actual distance between the two cities
Answer by fcabanski(1391)

1 inch : 20 miles is the ratio. Write it as a fraction: 1/20 (inches/miles).


The ratios are in proportion to each other (they are equal). inches/miles = 1/20 = 7/x


1/20 = 7/x : cross multiply and solve for x.


x=140 miles.